OpenStudy (beef123):

Use the rule y = -x + 7 to fill in the blank. If (-1, y) is on the graph, then y =

OpenStudy (will.h):

Substitute x = -1 in the expression And solve for y Y = -(-1) + 7 Y = 1 + 7 Y = 8 Make sense?

likeabossssssss (likeabossssssss):

\(\bbox[5pt,midnightblue]{\LARGE\color{yellow}{\star ^\bigstar Welcome~to~~os~ ^\bigstar \star}}\)

OpenStudy (beef123):

not really I don't think that that is right, but thanks anyway.